OpenStudy (anonymous):

1. Two planes intersect at _____ 2. What is the midpoint of coordinates 10 and -4? 3. What is the perimeter of a rectangle with width 6 inches and length 8 inches? 4. What is the area of a circle with a diameter of 16 ft. ? Use 3.14 for (pi) .

OpenStudy (anonymous):

1 is easy, just think of a coordinate plane. Where do the two axes connect, and what technically is that thing called? 2 Is vague, because to find the midpoint, you need two points, which 10 and -4 do not qualify as 3 To find the Perimeter, use the formula P=2w+2l 4 The area of a circles is (pi)r^2. 16=3.14r^2 16/3.14=r^2 Find the square root of sixteen divided by pi, and you'll have your radius